# latexfontFont shape

### v1

@@ 2.5. Links with HD automata
An automaton $\mathcal{A}$ is History-Deterministic (HD) if ond only if it is $1$-explorable, i.e. if there is a strategy $\sigma:\Sigma^{*}\to Q$ resolving the non-determinism based on the word read so far, with the guarantee that the run piloted by this strategy is accepting whenever the input word is in $L(\mathcal{A})$. See e.g. [BKKS13] for an introduction to HD automata.
@@ 2.5.1. Explorability in terms of HDness
2.5.2. Recognizing HD automata among explorable ones
In this section we give some motivation for considering explorable automata.
The arguments in this section are already hinted at in [BK18], and made explicit in the context of quantitative automata in [BL22]. We give them here for completeness, in order to provide some context for the relevance of the class of explorable automata.
One of the main open problems related to HD automata on infinite words is to decide, given a non-deterministic parity automaton, whether it is HD. For now, the problem is only known to be in PTime for coBüchi [KS15] and Büchi [BK18] automata. Extending this result even to $3$ parity ranks is still open, and only a naive ExpTime upper bound [HP06] is known in this case. The following result shows that explorability is relevant in this context:
Theorem 4.
Given an explorable parity automaton $\mathcal{A}$ of fixed parity index, it is in PTime to decide whether it is HD.
This is one of the motivations to get a better understanding of explorable automata. Indeed, if we can obtain an efficient algorithm for recognizing them, or if we are in a context guaranteeing that we are only dealing with explorable automata, this result shows that we can obtain an efficient algorithm for recognizing HD automata. Alternatively, even if membership to the class of explorable automata is provably hard to decide in general (as it will turn out), there can be some contexts where explorable automata are sufficient for the intended purposes. An example is given in [BL22], where it is shown that for LimSup and LimInf automata, Eve winning the game $G_{2}$ (defined below) implies that the automaton is explorable. Since Theorem 4 actually shows that Eve winning $G_{2}$ characterizes HDness for explorable automata, in this case it implies that the automaton is HD, as was shown in [BL22].
Let $\mathcal{A}$ be an explorable $[i,j]$-automaton.
We briefly recall the definition of the $k$-token game $G_{k}(\mathcal{A})$ defined in [BK18], for an arbitrary $k\in\mathbb{N}$. At each round, Adam plays a letter $a\in\Sigma$, then Eve moves her token according to an $a$-transition, and finally Adam moves his $k$ tokens according to $a$-transitions. Eve wins the play if her token builds an accepting run, or if all of Adam’s tokens build a rejecting run.
We will prove that the game $G_{2}(\mathcal{A})$ is won by Eve if and only $\mathcal{A}$ is HD.
First, it is clear that if $\mathcal{A}$ is HD, then Eve wins $G_{2}(\mathcal{A})$ [BK18]: Eve can simply play her HD strategy with her token, ignoring Adam’s tokens.
The interesting direction is the converse: we assume that Eve wins $G_{2}(\mathcal{A})$, and we show that under this assumption, $\mathcal{A}$ is necessarily HD. We use the following lemma:
Lemma 5 ([BK18, Thm. 14]).
Eve wins $G_{2}(\mathcal{A})$ if and only if Eve wins $G_{k}(\mathcal{A})$ for all $k\geq 2$.
Here we will combine this general result with the following lemma, that is specific to explorable automata:
Lemma 6.
If $\mathcal{A}$ is $k$-explorable and Eve wins $G_{k}(\mathcal{A})$, then $\mathcal{A}$ is HD.
Proof 2.3.
Let us note $Q=Q_{\mathcal{A}}$ the set of states of $\mathcal{A}$. We will build an explicit strategy witnessing that $\mathcal{A}$ is HD.
Let $\tau_{k}$ be a winning strategy for Determiniser in the $k$-explorability game of $\mathcal{A}$, and $\sigma_{k}$ a winning strategy for Eve in $G_{k}(\mathcal{A})$.
Let us explicit in detail the shape of these strategies. The strategy $\tau_{k}$ has access to the history of the play in the $k$-explorability game, and must decide on a move for Determiniser. Notice that in absence of memory limitations, it is always enough to know the history of the opponent’s moves (here the letters of $\Sigma$ played so far), since this allows to compute the answer of Determiniser at each step, and therefore build a unique play. Thus, we can take for $\tau_{k}$ a function $\Sigma^{*}\to Q^{k}$. If the word played so far is $u\in\Sigma^{*}$, the tuple of states reached by the $k$ tokens moved according to $\tau_{k}$ is $\tau_{k}(u)\in Q^{k}$. In particular $\tau_{k}(\varepsilon)=(q_{0}^{\mathcal{A}},\dots,q_{0}^{\mathcal{A}})$.
If $w=a_{1}a_{2}\dots\in\Sigma^{\omega}$, and $i\in\mathbb{N}$, let us note $(q_{w,1}^{i},\dots,q_{w,k}^{i})=\tau_{k}(a_{1}\dots a_{i})$. That is $q_{w,j}^{i}$ is the state reached by the $j^{\mathrm{th}}$ token after $i$ steps in the run induced by $\tau_{k}$ and $w$. If $j\in[1,k]$, let us note $\rho_{w,j}$ the infinite run $q_{w,j}^{0}q_{w,j}^{1}q_{w,j}^{2}\dots$, followed by the $j^{\mathrm{th}}$ token in this play. Since $\tau_{k}$ is a winning strategy in the $k$-explorability game of $\mathcal{A}$, we have the guarantee that for all $w\in L(\mathcal{A})$, there exists $j\in[1,k]$ such that $\rho_{w,j}$ is accepting.
If $u=a_{1}\dots a_{n}\in\Sigma^{*}$ is a finite word, we define $\tau_{k}^{\prime}(u)=(\tau_{k}(\varepsilon),\tau_{k}(a_{1}),\tau_{k}(a_{1}a_{2})\dots,\tau_{k}(u))$, this is a description of the $k$ partial runs induced by $\tau_{k}$ on $u$.
Let us now turn to the strategy $\sigma_{k}$ of Eve in $G_{k}(\mathcal{A})$. The type of this strategy is $\sigma_{k}:\Sigma^{*}\times(Q^{k})^{*}\to Q$. Indeed, this time, the history of Adam’s moves must contain his choice of letters together with his choices of positions for his $k$ tokens. So $\sigma_{k}(u,\gamma)$ gives the state reached by Eve’s unique token after a history $(u,\gamma)$ for the moves of Adam. Notice that at each step, Eve must move before Adam in this game $G_{k}(\mathcal{A})$, so $\gamma$ does not contain the choice of Adam on the last letter of $u$. This means that except for $u=\varepsilon$, we can always assume $|u|=|\gamma|+1$ in a history $(u,\gamma)$.
We have the guarantee that if Adam plays an infinite word $w$ together with runs $\rho_{1},\dots,\rho_{k}$ on $w$, such that at least one of these runs is accepting, then the run yielded by $\sigma_{k}$ against $(w,(\rho_{1},\dots,\rho_{k}))$ is accepting.
We finally define the HD strategy $\sigma$ for $\mathcal{A}$, of type $\Sigma^{*}\to Q$, by induction: $\sigma(\varepsilon)=q_{0}^{\mathcal{A}}$, and $\sigma(ua)=\sigma_{k}(ua,\tau_{k}^{\prime}(u))$.
This amounts to playing the strategy $\sigma_{k}$ in $G_{k}(\mathcal{A})$, against Adam playing a word $w$ and moving his $k$ tokens according to the strategy $\tau_{k}$ against $w$. If the infinite word $w=a_{1}a_{2}\dots$ chosen by Adam is in $L(\mathcal{A})$, then by correctness of $\tau_{k}$ one of the $k$ runs $\rho_{w,1},\dots,\rho_{w,k}$ yielded by $\tau_{k}$ is accepting. Hence, by correctness of $\sigma_{k}$, the run $\sigma(\varepsilon)\sigma(a_{1})\sigma(a_{1}a_{2})$ yielded by $\sigma$ (based on $\sigma_{k}$) is accepting. This concludes the proof that $\sigma$ is a correct HD strategy for $\mathcal{A}$, witnessing that $\mathcal{A}$ is HD.
Combining the results of this section, we obtain that if $\mathcal{A}$ is explorable, then $\mathcal{A}$ is HD if and only if Eve wins $G_{2}(\mathcal{A})$. It remains to show that the winner of $G_{2}(\mathcal{A})$ can be computed in PTime, for fixed parity index of $\mathcal{A}$. This is already stated without proof in the conclusion of [BK18], but let us explicit it here for completeness, thereby achieving the proof of Theorem 4.
Lemma 7.
For parity automata of fixed parity index $[i,j]$, it is in PTime to decide the winner of $G_{2}(\mathcal{A})$.
Proof 2.4.
[…]
  • From a position $((p,q_{1},q_{2},a,E),q_{\mathcal{D}},\_)$, when Eve chooses a successor position in $G_{2}(\mathcal{A})$ according to a transition $\delta=(p,a,p^{\prime})$, the game $G^{\prime}$ will move to $((p^{\prime},q_{1},q_{2},a,A),q_{\mathcal{D}},\mathrm{rk}(\delta))$, i.e. remembering in the last component the parity rank seen by Eve’s token.
  • From a position $((p^{\prime},q_{1},q_{2},a,A),q_{\mathcal{D}},k)$, when Adam chooses a successor position in $G_{2}(\mathcal{A})$ according to transitions $\delta_{1}=(q_{1},a,q_{1}^{\prime})$ and $\delta_{2}=(q_{2},a,q_{2}^{\prime})$, the game $G^{\prime}$ will move to $((p^{\prime},q_{1}^{\prime},q_{2}^{\prime}),\delta_{\mathcal{D}}(q_{\mathcal{D}},(k,\mathrm{rk}(\delta_{1}),\mathrm{rk}(\delta_{2}))),\_)$, i.e. the automaton $\mathcal{D}$ will deterministically advance according to the parity ranks seen on the $3$ tokens.
  • Eve’s winning condition in $G^{\prime}$ is the parity condition of $\mathcal{D}$.
Such a composition of a game with a deterministic automaton is standard (up to the bookkeeping needed here to remember Eve’s transition rank), and we obtain that $G^{\prime}$ and $G_{2}(\mathcal{A})$ have same winner: the winning condition of $G_{2}(\mathcal{A})$ is simply taken care of by $\mathcal{D}$ in $G^{\prime}$, allowing to simplify the winning condition of the game. Since $\mathcal{D}$ is of fixed size, $G^{\prime}$ is still of polynomial size in $\mathcal{A}$, and it is a parity game, so we can decide its winner in PTime. This completes the description of the PTime algorithm algorithm to decide the winner of $G_{2}(\mathcal{A})$: build the game $G^{\prime}$ by composing $G_{2}(\mathcal{A})$ with a fixed deterministic parity automaton $\mathcal{D}$, and solve the game $G^{\prime}$.
@@ 3. Expressivity of explorable automata
We will now generalize this to $[0,2]$-automata:
@@
Proof 3.3.
Just as in the above proof, the converse inclusion is straightforward, as a deterministic automaton is always explorable. Let $L$ be a language recognized by some $k$-explorable $[0,2]$-automaton $\mathcal{A}$. Similarly as above, this means that the union product $\mathcal{A}^{k}$ is HD, and this is witnessed by a HD strategy with finite memory $M$. This allows us to build a deterministic automaton $\mathcal{B}$ with states $Q^{k}\times M$, where the acceptance condition only depends on the $Q^{k}$ component, and is a union of $k$ $[0,2]$-conditions. In order to obtain a deterministic $[0,2]$-automaton, we need to compose $\mathcal{B}$ with a deterministic $[0,2]$-automaton $\mathcal{C}$ on alphabet $\Gamma:=[0,2]^{k}$, that accepts an infinite word if and only if one of its $k$-components is $[0,2]$-accepting. This can be obtained, by combining the well-known breakpoint construction on the $[0,1]$ part, with the construction from Lemma 9, with ranks $2$ playing the role of Büchi states.
Intuitively, the automaton will remember which components has seen a $1$ since the last reset. Seeing a $2$ anywhere causes to produce a $2$ globally, and reset the memory. On intervals without any $2$, if a $1$ has been seen on every component, the automaton produces a $1$ globally and resets its memory. The automaton $\mathcal{C}=(\Gamma,Q_{\mathcal{C}},q_{0}^{\mathcal{C}},\delta_{\mathcal{C}})$ can be built as follows:
  • $Q_{\mathcal{C}}=\{0,1\}^{k}$
  • $q_{0}^{\mathcal{C}}=(0,0,\dots,0)$
@@
    $\delta_{\mathcal{C}}((a_{1},\dots,a_{k}),(b_{1},\dots,b_{k}))=\left\{\begin{array}[]{llr}(q_{0}^{\mathcal{C}},2)&\text{ if some $b_{i}$ is $2$,}\\ (q_{0}^{\mathcal{C}},1)&\text{ otherwise if for all $i$, $a_{1}=1$ or $b_{i}=1$,}\\ ((a_{i}\vee b_{i})_{1\leq i\leq k},0)&\text{ otherwise.}\end{array}\right.$
@@
Similarly to what was done in Lemma 7, it now suffices to compose $\mathcal{B}$ with $\mathcal{C}$, i.e. having $\mathcal{C}$ read the ranks output by $\mathcal{B}$, and using the $[0,2]$ acceptance condition of $\mathcal{C}$. This yields a deterministic $[0,2]$-automaton recognizing $L$, represented in Figure 3.
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$a:1$
$\mathcal{A}$
$\mathcal{A}$
$\mathcal{A}$
$1$
00
00
$1$
${b:0}$
00
$1$
00
00
$1$
00
$1$
$1$
00
00
00
$2$
$2$
$1$
$b:2$
00
00
00
$M$
$M$
$M$
$M$
Figure 3. The composition of $\mathcal{C}$ with $\mathcal{B}=\mathcal{A}^{k}\times M$ (here with $k=3$). Internal ranks/memory states labelled $1$ are represented in red, until the reset to $q_{0}^{\mathcal{C}}$ (in green) is performed when a $1$ is reached on each component. Input letters and global ranks of the resulting automaton are shown on top.
@@
For $l$ even, we define $\mathcal{A}_{l}$ as the copy of $\mathcal{A}$ where all priorities $<l$ are replaced with $1$, all priorities $>l$ are replaced with $3$, and all priorities equal to $l$ are replaced with $2$. $\mathcal{A}_{l}$ is thus a deterministic $[1,3]$-automaton. It has the notable property that a word $w$ is accepted by $\mathcal{A}_{l}$ if and only if it is accepted by $\mathcal{A}$ with highest priority $l$. Therefore $\mathcal{A}$ is equivalent to the union of the $\{\mathcal{A}_{l}~{}|~{}l\text{ even }\in[1,d]\}$. We thus build the automaton $\mathcal{A}^{\prime}$ where the initial state branches non-deterministically via $\varepsilon$-transitions towards all the different $\mathcal{A}_{l}$ for all even $l\in[1,d]$. This automaton is non-deterministic, of parity index $[1,3]$, and recognizes exactly the words recognized by $\mathcal{A}$. It is $\frac{d}{2}$-explorable, as this is the maximum number of tokens needed to place one in each $\mathcal{A}_{l}$ at the start, after which their progression becomes deterministic.
This construction can actually by applied to non-deterministic automata as well, yielding the following result:
@@ 4. Decidability and complexity of the explorability problem
Let us start by recalling the population control problem (PCP) of [BDG+19]. {defi}[$k$-population game] Given an NFA $\mathcal{B}$ with a distinguished target state $f\in Q_{\mathcal{B}}$, and an integer $k\in\mathbb{N}$, the $k$-population game is played similarly to the $k$-explorability game, only the winning condition differs: Spoiler wins if the game reaches a position where all tokens are in the state $f$.
The PCP asks, given $\mathcal{B}$ and $f\in Q_{\mathcal{B}}$, whether Spoiler wins the $k$-population game for all $k\in\mathbb{N}$. Notice that this convention is opposite to explorability, where positive instances are defined via a win of Determiniser. The PCP is shown in [BDG+19] to be ExpTime-complete. We will present here a direct exponential reduction from the explorability problem to the PCP.
Theorem 14 (Direct reduction to the PCP).
@@
Let $\mathcal{A}=(\Sigma,Q_{\mathcal{A}},q_{0}^{\mathcal{A}},F_{\mathcal{A}},\Delta_{\mathcal{A}})$ be an NFA. Our goal is to build an exponential NFA $\mathcal{B}$ with a distinguished state $f$ such that $(\mathcal{B},f)$ is a negative instance of the PCP if and only if $\mathcal{A}$ is explorable.
@@
The initial state of $\mathcal{B}$ is $q_{0}^{\mathcal{B}}=(q_{0}^{\mathcal{A}},\{q_{0}^{\mathcal{A}}\})$. Notice that we do not need to specify accepting states or priorities in $\mathcal{B}$, as acceptance plays no role in the PCP.
@@
[…]
@@
Notice that as long as letters of $\Sigma$ are played, the second component of states of $\mathcal{B}$ evolves deterministically and keeps track of the set of reachable states in $\mathcal{A}$. Moreover, the letter $a_{\mathrm{test}}$ also acts deterministically on $Q_{\mathcal{B}}$. Therefore, the only non-determinism to be resolved in $\mathcal{B}$ is how the first component evolves, which amounts to building a run in $\mathcal{A}$. Thus, strategies driving tokens in $\mathcal{A}$ and $\mathcal{B}$ are isomorphic. It now suffices to observe that Spoiler wins the $k$-population game on $(\mathcal{B},f)$ if and only if he has a strategy allowing to eventually play $a_{\mathrm{test}}$ while all tokens are in a state of the form $(q,X)$ with $q\notin F_{\mathcal{A}}$ and $X\cap F_{\mathcal{A}}\neq\emptyset$. This is equivalent to Spoiler winning the $k$-explorability game of $\mathcal{A}$, since $X\cap F_{\mathcal{A}}\neq\emptyset$ witnesses that the word played so far is in $L(\mathcal{A})$.
This concludes the proof that $\mathcal{A}$ is explorable if and only if $(\mathcal{B},f)$ is a negative instance of the PCP. So given an NFA $\mathcal{A}$ that we want to test for explorability, it suffices to build $(\mathcal{B},f)$ as above, and use the ExpTime algorithm from [BDG+19] as a black box on $(\mathcal{B},f)$. Since $\mathcal{B}$ is of exponential size compared to $\mathcal{A}$, this achieves the proof of Theorem 14.

### v2

@@ 2.5. Links with HD automata
An automaton $\mathcal{A}$ is History-Deterministic (HD) if and only if it is $1$-explorable, i.e. if there is a strategy $\sigma:\Sigma^{*}\to Q$ resolving the non-determinism based on the word read so far, with the guarantee that the run piloted by this strategy is accepting whenever the input word is in $L(\mathcal{A})$. See e.g. [BKKS13] for an introduction to HD automata.
@@ 3. Expressivity of explorable automata
Let us first state a general lemma on expressivity of explorable automata with any $\omega$-regular acceptance condition.:
Lemma 7.
Let $\mathit{Acc}$ be any $\omega$-regular acceptance condition. Then a language $L$ is recognized by an explorable $\mathit{Acc}$-automaton if and only if $L$ is a finite union of languages recognized by deterministic $\mathit{Acc}$-automata.
Proof 3.3.
Let us start with the right-to-left direction. Let $L=\bigcup_{1\leq i\leq k}L_{i}$ , where each $L_{i}$ is recognized by a deterministic $\mathit{Acc}$-automaton $\mathcal{A}_{i}$. We can build an explorable $\mathit{Acc}$-automaton $\mathcal{A}$ recognizing $L$ by having its initial state branch non-deterministically towards each $\mathcal{A}_{i}$. This automaton is clearly $k$-explorable, and recognizes $L$.
Assume now that $L$ is recognized by some $k$-explorable $\mathit{Acc}$-automaton $\mathcal{A}$ with state space $Q$. Let $\sigma$ be a winning strategy in the $k$-token game of $\mathcal{A}$, with finite memory $M$. The existence of such a strategy is guaranteed by the fact that $\omega$-regular languages are closed under union, and any $\omega$-regular game is finite-memory determined. Composing $\mathcal{A}$ with $\sigma$ yields a deterministic automaton $\mathcal{B}$ of state space $Q^{k}\times M$. Let $\mathcal{A}_{i}$ be the automaton obtained from $\mathcal{B}$ by considering the $\mathit{Acc}$-acceptance condition of the $i$-th component. Each $\mathcal{A}_{i}$ is a deterministic $\mathit{Acc}$-automaton, and we have $L=\bigcup_{1\leq i\leq k}L(\mathcal{A}_{i})$.
The results obtained in the rest of the section are actually corollaries of Lemma 7, by instantiating $\mathit{Acc}$ to the different parity conditions: the $[0,2]$ condition is closed under union (yielding Lemma 8), and any parity condition is a union of $[1,3]$ conditions (yielding Theorem 9). However, we will still provide direct proofs for better efficiency of the constructions. Indeed, in the above lemma, each component needs to incorporate a memory structure updating all $k$ tokens, which can be avoided in the direct proofs below.
Let us treat the case of $[0,2]$-automata:
@@
Proof 3.4.
Just as in the proof of Lemma 5, the converse inclusion is straightforward, as a deterministic automaton is always explorable. Let $L$ be a language recognized by some $k$-explorable $[0,2]$-automaton $\mathcal{A}$. Similarly as above, this means that the union product $\mathcal{A}^{k}$ is HD, and this is witnessed by an HD strategy with finite memory $M$. This allows us to build a deterministic automaton $\mathcal{B}$ with states $Q^{k}\times M$, where the acceptance condition only depends on the $Q^{k}$ component, and is a union of $k$ $[0,2]$-conditions. In order to obtain a deterministic $[0,2]$-automaton, we need to compose $\mathcal{B}$ with a deterministic $[0,2]$-automaton $\mathcal{C}$ on alphabet $\Gamma:=[0,2]^{k}$, that accepts an infinite word if and only if one of its $k$-components is $[0,2]$-accepting.
Intuitively, the automaton will cycle through the components. Seeing a $1$ in the current component will prompt the automaton to go to the next component while producing a $1$ globally. Seeing a $2$ anywhere causes to produce a $2$ globally (and reset to the first component). The automaton $\mathcal{C}=(\Gamma,Q_{\mathcal{C}},q_{0}^{\mathcal{C}},\delta_{\mathcal{C}})$ can be built as follows:
  • $Q_{\mathcal{C}}=\{1,\dots,k\}$
  • $q_{0}^{\mathcal{C}}=1$
@@
[…]
@@
It now suffices to compose $\mathcal{B}$ with $\mathcal{C}$, i.e. having $\mathcal{C}$ read the ranks output by $\mathcal{B}$, and using the $[0,2]$ acceptance condition of $\mathcal{C}$. This yields a deterministic $[0,2]$-automaton with state space $Q_{\mathcal{B}}\times\mathcal{Q}_{\mathcal{C}}$, recognizing $L$.
@@
For $l$ even, we define $\mathcal{A}_{l}$ as the copy of $\mathcal{A}$ where all priorities <l<l are replaced with $1$, all priorities >l>l are replaced with $3$, and all priorities equal to $l$ are replaced with $2$. $\mathcal{A}_{l}$ is thus a deterministic $[1,3]$-automaton. It has the notable property that a word $w$ is accepted by $\mathcal{A}_{l}$ if and only if it is accepted by $\mathcal{A}$ with highest priority $l$. Therefore, $\mathcal{A}$ is equivalent to the union of the $\{\mathcal{A}_{l}~|~l\text{ even }\in[1,d]\}$. We thus build the automaton $\mathcal{A}^{\prime}$ where the initial state branches non-deterministically via $\varepsilon$-transitions towards all the different $\mathcal{A}_{l}$ for all even $l\in[1,d]$. This automaton is non-deterministic, of parity index $[1,3]$, and recognizes exactly the words recognized by $\mathcal{A}$. It is $\frac{d}{2}$-explorable, as this is the maximum number of tokens needed to place one in each $\mathcal{A}_{l}$ at the start, after which their progression becomes deterministic.
This construction can actually be applied to non-deterministic automata as well, yielding the following result:
@@ 4. Decidability and complexity of the explorability problem
Let us start by recalling the population control problem of [BDG+19]. {defi}[$k$-population game] Given an NFA $\mathcal{B}$ with a distinguished state $f\in Q_{\mathcal{B}}$, and an integer $k\in\mathbb{N}$, the $k$-population game is played similarly to the $k$-explorability game, only the winning condition differs: Spoiler wins if the game reaches a position where all tokens are in the state $f$.
The population control problem asks, given $\mathcal{B}$ and $f\in Q_{\mathcal{B}}$, whether Spoiler wins the $k$-population game for all $k\in\mathbb{N}$. Notice that this convention is opposite to explorability, where positive instances are defined via a win of Determiniser. The population control problem is shown in [BDG+19] to be ExpTime-complete. We will present here a direct exponential reduction from the explorability problem to the population control problem.
Theorem 11 (Direct reduction to the population control problem).
@@
Let $\mathcal{A}=(\Sigma,Q_{\mathcal{A}},q_{0}^{\mathcal{A}},F_{\mathcal{A}},\Delta_{\mathcal{A}})$ be an NFA. Our goal is to build an exponential NFA $\mathcal{B}$ with a distinguished state $f$ such that $(\mathcal{B},f)$ is a negative instance of the population control problem (i.e. the player controlling the tokens can prevent their accumulation in $f$) if and only if $\mathcal{A}$ is explorable.
@@
The initial state of $\mathcal{B}$ is $q_{0}^{\mathcal{B}}=(q_{0}^{\mathcal{A}},\{q_{0}^{\mathcal{A}}\})$. Notice that we do not need to specify accepting states or priorities in $\mathcal{B}$, as acceptance plays no role in the population control problem.
@@
  • $(p,X)\overset{a}{\longrightarrow}(q,\Delta_{\mathcal{A}}(X,a))$ if $a\in\Sigma$ and $q\in\Delta_{\mathcal{A}}(p,a)$,
@@
Proof 4.1.
Notice that as long as letters of $\Sigma$ are played, the second component of states of $\mathcal{B}$ evolves deterministically and keeps track of the set of reachable states in $\mathcal{A}$. Moreover, the letter $a_{\mathrm{test}}$ also acts deterministically on $Q_{\mathcal{B}}$. Therefore, the only non-determinism to be resolved in $\mathcal{B}$ is how the first component evolves, which amounts to building a run in $\mathcal{A}$. Thus, strategies driving tokens in $\mathcal{A}$ and $\mathcal{B}$ are isomorphic. It now suffices to observe that Spoiler wins the $k$-population game on $(\mathcal{B},f)$ if and only if he has a strategy allowing to eventually play $a_{\mathrm{test}}$ while all tokens are in a state of the form $(q,X)$ with $q\notin F_{\mathcal{A}}$ and $X\cap F_{\mathcal{A}}\neq\emptyset$. This is equivalent to Spoiler winning the $k$-explorability game of $\mathcal{A}$, since $X\cap F_{\mathcal{A}}\neq\emptyset$ witnesses that the word played so far is in $L(\mathcal{A})$. This concludes the proof that $\mathcal{A}$ is explorable if and only if $(\mathcal{B},f)$ is a negative instance of the population control problem.
Given an NFA $\mathcal{A}$ that we want to test for explorability, it suffices to build $(\mathcal{B},f)$ as above, and use the ExpTime algorithm from [BDG+19] as a black box on $(\mathcal{B},f)$. Since $\mathcal{B}$ is of exponential size compared to $\mathcal{A}$, this achieves the proof of Theorem 11.